find the area of a triangle with points A(-6,-10), B(4,-7), and C(-1,3)
Answer by Alan3354(69443)

find the area of a triangle with points A(-6,-10), B(4,-7), and C(-1,3)
==============================
Arrange the points like this:
---
..A....B....C....A
-6.....4...-1...-6
-10...-7....3..-10
----
Add the diagonal products starting at the upper left.
-6*-7 + 4*3 + -1*-10 = 42 + 12 + 10 = 64
---
Add the diagonal products starting at the lower left.
-10*4 + -7*-1 + 3*-6 = -40 + 7 - 18 = -51
--------
The difference = 125
The area = 1/2 that, = 62.5 sq units